A particle is performing SHM along x axis with amplitude 4cm and time period 1.2seconds.The minimum time taken by particle to move from x=2cm to=4cm and back again is given by

Grade:12

1 Answers

Khimraj
6 years ago
l time period T = 1.2 sec
Time taken by to particle to move from x=2cm to x=4cm will be T/6 and to come back will aslo be T/6.
So total time taken will be equal t0 T/6 + T/6 = T/3 = 0.4 sec.
Hope it clears.